# Beyond Training: Dynamic Token Merging for Zero-Shot Video Understanding

### v1

@@ Abstract
Recent advancements in multimodal large language models (MLLMs) have opened new avenues for video understanding. However, achieving high fidelity in zero-shot video tasks remains challenging. Traditional video processing methods rely heavily on fine-tuning to capture nuanced spatial-temporal details, which incurs significant data and computation costs. In contrast, training-free approaches, though efficient, often lack robustness in preserving context-rich features across complex video content. To this end, we propose DyTo, a novel dynamic token merging framework for zero-shot video understanding that adaptively optimizes token efficiency while preserving crucial scene details. DyTo integrates a hierarchical frame selection and a bipartite token merging strategy to dynamically cluster key frames and selectively compress token sequences, striking a balance between computational efficiency with semantic richness. Extensive experiments across multiple benchmarks demonstrate the effectiveness of DyTo, achieving superior performance compared to both fine-tuned and training-free methods and setting a new state-of-the-art for zero-shot video understanding.
@@ 1 Introduction
Figure 1: Comparison with baseline Video LLMs under 7B model size on 10 video benchmarks. DyTo and other baselines are marked using solid (—) and dashed (- - -) lines, respectively. DyTo outperforms existing training-free methods on almost benchmarks and achieves even better results than most SFT methods fine-tuned on video datasets.
@@ 2.1 Video Large Language Models
Video language model has witnessed significant advancements over the past years [18, 7, 40, 28, 34]. Video-ChatGPT [1] retrieves features from each frame and subsequently combines them using two operations for spatial and temporal pooling prior to feeding them into a large language model. LLaMA-VID [17] utilized a dual token approach to effectively compress the video token by differentiating between context and content. Vista-LLaMA [21] presented EDVT-Attention along with a sequential vision projector that emphasizes visual tokens while decreasing temporal tokens by sequentially merging them using a Q-former. Video-LLaVA [18] aligns the encoders for images and videos beforehand, enabling shared projections and joint training across both image and video tasks, thus mapping them into the language space. VideoChat [15] utilized cross-attention to condense video tokens alongside user inquiries and conversational context. Video-LLaMA [38] introduces a Video Q-Former and an Audio Q-Former, allowing for the integration of multiple modalities in video understanding. In contrast, Video-LLaMA2 [7] designs a spatial-temporal convolution connector to replace the Q-Former for spatial-temporal representation learning. Chat-UniVi [10] developed a unified model for images and videos that uses dynamic token merging with k-NN to simplify spatial and temporal tokens. SeViLA [36] concentrated on detecting and extracting keyframes relevant to the inquiries and analyzed the video by transforming these keyframes into video tokens.
@@ 3 Method
Figure 2: Overview of DyTo, which is a training-free model built upon Image MLLM without continued fine-tuning. Our method utilizes the [CLS] token(pink block) to divide the video into K clusters, then the dynamic bipartite merging module samples the frames from every cluster and controls the final output length as $L_{out}$.
@@ 3.1 Coarse-grained Hierarchical Clustering
The 1-NN temporal graph $G$ defines an adjacency matrix where each node is connected to its nearest neighbor based on the temporally weighted distances $W_{t}$. We make the links symmetric by setting $G(j,i)=1$ to encode space-temporal distances and form connected components to clusters conveniently. The connected components of Graph in Equ.2 automatically partition the data into discovered clusters. We then merge these clusters recursively based on the cluster averages of features when only one cluster remains and select the second-largest cluster $K$ for our video segmentation. We denote the set of indexes of the frames in the cluster as $K_{n}$. Therefore, the set of frames within the $n_{th}$ cluster $K_{n}$ can be formulated as:
$V_{n}^{\prime}=\{V_{i}|i\in K_{n}\}$ (3)
We uniformly sample frames from every cluster and combine them as the keyframe sequence $V_{k}=\{V^{\prime}_{1},V^{\prime}_{2},...,V^{\prime}_{K}\}\in\mathbb{R}^{(L-1)\times D}$ from $K$ segmentations. The segmentation can yield the distribution of events contained in the current video sequence, thereby guiding the fine-grained merging in subsequent steps.
@@ 3.2 Fine-grained Dynamic Bipartite Merging
Previous method [12, 35, 31] either simply concatenates frames and designs prompts for the VLM or uses pooling method on each image token to construct the video representation. DyTo propose a novel method to effectively construct video representations from selected frames. After acquiring the segmentations of the frame sequence through the previously described method, we implement a dynamic bipartite token merging approach inspired by ToMe [3] to minimize the number of visual tokens. For each individual frame in $V_{k}$, we sequentially divide the $R_{i}$ tokens into two non-overlap tokens sets $\mathbb{P}$ with $r_{i}$ tokens and $\mathbb{Q}$ with $R_{i}-r_{i}$ tokens at the $i^{th}$ steps, where initial $R_{0}=L-1$. We set the $r_{i}=R_{0}-\frac{Z}{K}$ to dynamically merge the image to preserve more tokens under the fixed visual token length $Z$. To obtain the similarity scores, each visual token is split into $H$ heads along channel dimensions, each with $D/H$ channels. The similarity score for each token pair is obtained by averaging the cosine similarity scores over all heads following Eq. (4)
@@ 4.1 Evaluation Benchmarks
For structured VQA, we evaluate DyTo on a diverse set of multiple-choice benchmarks, including NExTQA [32], VideoMME [8], EgoSchema [24], STAR [30], and IntentQA [13], all of which are designed to quantify the model’s accuracy on selecting the correct answer among predefined options. Collectively, these benchmarks provide a comprehensive evaluation of DyTo to interpret complex, multimodal data and to produce accurate, contextually rich responses across varying levels of task structure. In addition, we perform experiments on the VideoMME [8] benchmark under the “w/o subs” configuration, which restricts access to subtitles, thereby isolating the model’s reliance on visual and temporal cues in video understanding.
@@
For open-ended VQA tasks, we assess the zero-shot performance of DyTo on MSVD-QA [4], MSRVTT-QA [33], TGIF-QA [16], and ActivityNet-QA [37] benchmarks. These benchmarks require the model to autonomously generate free-form responses, simulating real-world question-answering contexts that demand nuanced understanding of video content. Consistent with prior methodologies, we employ the Video-Chatgpt Generation benchmarks[22](VCG Benchmark) as a principal benchmark for evaluating text generation capabilities, assessed across five key dimensions, including Correctness of Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU), and Consistency (CO). Following Wu [31], we utilize GPT-3.5-Turbo-0125 to ensure fair comparisons with other methods.
@@ 4.2 Experimental Setup
Implementation details.
All evaluations could be conducted on a single Nvidia A100 80G graphics card. To accelerate inference, we use a Linux server equipped with 8 Nvidia A100 80G cards. DyTo is based on the LLaVA-NeXT model series, with pre-trained weights available on HuggingFace [^1] https://huggingface.co/collections/liuhaotian/llava-16-65b9e40155f60fd046a5ccf2. We implement rotary position embedding (RoPE) and apply a scaling factor of 2, extending the original context length from 4096 to 8192 tokens.
@@
In our approach, we uniformly sample each video to $N=100$ frames. Each frame is resized to $336\times 336$ to match the input dimensions of the visual encoder, which then outputs $24\times 24$+[CLS] visual tokens plus a [CLS] token. Following Sarfraz et al. [27], we collect the [CLS] token as the coarse-grained feature for each frame and subsequently partition the sequence into $K$ clusters, using $N$ iterations of hierarchical clustering3.1) to group semantically similar frames.
Due to hardware limitations, we set the visual token sequence length $L_{out}$ to either 3680 or 7200, corresponding to model sizes of 7B and 34B, respectively. And to optimize feature representation while controlling token redundancy, we dynamically adjust the merge ratio using $r=L/K$. Additionally, we set the number of heads $H=16$, aligning with CLIP-VIT/L. Following Wu [31], Xu et al. [35], we fix the initial merge ratio $r_{1}=288$ in the first iteration to ensure optimal performance across the sequence.
@@ 4.3 Main Results
Method LLM Video Vision NExTQA EgoSchema IntentQA STAR VideoMME
Size Trained Encoder acc. acc. acc. acc. acc.
InternVideo 7B CLIP-L O 59.1 32.1 - 41.6 -
VideoChat2 7B UMT-L O 61.7 - - 59.0 -
Sevilla 7B CLIP-L O 63.6 - 60.9 44.6 -
Video-LLaVA 7B ViT-L O - - - - 39.9
Chat-UniVi-v1.5 7B CLIP-G O - - - - 40.6
DeepStack-L 7B CLIP-L X 61.0 38.4 - - -
IG-VLM 7B CLIP-L X 63.1 35.8 60.1 49.6 39.8
SlowFast-LLaVA 7B CLIP-L X 64.2 45.5 60.1 49.0 40.4
DyTo 7B CLIP-L X 65.7 48.6 61.6 50.7 41.2
Table 1: DyTo performance on multiple-choice benchmarks compared to leading fine-tuned and training-free models. Bold and underline indicate best performance with and without methods that require additional fine-tuning. Results highlight the superior zero-shot performance of DyTo across various datasets, especially for tasks requiring detailed temporal and contextual comprehension.
As shown in Table 1, DyTo consistently outperforms several fine-tuned and training-free models across various benchmarks, including ExTQA [32], VideoMME [8], EgoSchema [24], STAR [30], and IntentQA [13]. Notably, DyTo demonstrated its adaptability to diverse video understanding contexts, achieving superior accuracy even compared to models that employ extensive fine-tuning. For example, on NExTQA, DyTo achieved an accuracy of 65.7%, setting a new state-of-the-art performance. Similarly, performance on EgoSchema, IntentQA, STAR, and VideoMME illustrate the effectiveness of DyTo in handling task-specific reasoning within video content, often by maintaining accuracy advantages in scenarios requiring nuanced temporal and contextual understanding.
Our experiments indicate that the hierarchical clustering and bipartite merging strategies contribute significantly to its accuracy. The clustering method captures critical events across frames while preserving semantic richness, supporting robust question-answering performance without incurring high computational overhead. Moreover, the proposed dynamic token merging better retained crucial contextual information than traditional pooling methods, which may overly simplify token representations, potentially leading to loss of action details critical for accurate response selection.
@@
Method LLM Vision Video MSVD-QA MSRVTT-QA TGIF-QA ANet-QA Video-ChatGPT (VCG Benchmark)
Size Encoder Trained acc./score acc./score acc./score acc./score CI DO CU TU CO Average
Video-LLaMA 7B CLIP-G O 51.6/2.5 29.6/1.8 - 12.4/1.1 1.96 2.18 2.16 1.82 1.79 1.98
Video-LLama2 7B CLIP-L O 70.9/3.8 - - 50.2/3.3 3.16 3.08 3.69 2.56 3.14 3.13
Video-ChatGPT 7B CLIP-L O 64.9/3.5 49.3/2.9 51.4/3.0 35.2/2.7 2.50 2.57 2.69 2.16 2.20 2.42
VideoGPT+ 3.8B CLIP-L O 72.4/3.9 60.6/3.6 74.6/4.1 50.6/3.6 3.27 3.18 3.74 2.83 3.39 3.28
Video-LLava 7B ViT-L O 70.7/3.9 59.2/3.5 70.0/4.0 45.3/3.3 - - - - - -
MovieChat 7B CLIP-G O 75.2/3.8 52.7/2.6 - 45.7/3.4 2.76 2.93 3.01 2.24 2.42 2.67
MovieChat+ 7B CLIP-G O 76.5/3.9 53.9/2.7 - 48.1/3.4 - - - - - -
VideoChat2 7B UMT-L O 70.0/3.9 54.1/3.3 - 49.1/3.3 3.02 2.88 3.51 2.26 2.81 2.98
Vista-LLAMA 7B CLIP-G O 65.3/3.6 60.5/3.3 - 48.3/3.3 2.44 2.64 3.18 2.26 2.31 2.57
LLama-VID 13B CLIP-L O 69.7/3.7 57.7/3.2 - 47.4/3.3 2.96 3.00 3.53 2.46 2.51 2.89
PLLaVA 7B CLIP-L O 76.6/4.1 62.0/3.5 77.5/4.1 56.3/3.5 - - - - - -
LLaVA-NeXT-Video 7B CLIP-L O - - - 53.5/3.5 - - - - - -
FreeVA 7B CLIP-L X 71.2/3.8 58.5/3.2 - 53.7/3.5 - - - - - -
IG-VLM 7B CLIP-L X 78.3/3.9 63.7/3.4 72.7/4.0 53.8/3.2 3.11 2.78 3.48 2.44 3.29 3.03
SlowFast-LLaVA 7B CLIP-L X 78.7/3.9 66.2/3.4 77.5/4.0 53.9/3.1 3.09 2.70 3.57 2.52 3.35 3.04
DyTo 7B CLIP-L X 77.6/3.9 64.1/3.4 78.0/4.0 54.3/3.2 3.12 2.93 3.72 2.52 3.52 3.16
@@
In open-ended settings, as shown in Table 2, DyTo demonstrates competitive zero-shot capabilities. It consistently outperforms existing methods on benchmarks including MSVD-QA [4], MSRVTT-QA [33], TGIF-QA [16], and ActivityNet-QA [37]. Additionally, DyTo performs exceptionally well on Video-ChatGPT [22], surpassing even those methods that require fine-tuning, highlighting its robustness in zero-shot settings. DyTo ’s ability to maintain high accuracy without fine-tuning underscores the strength of its adaptive framework in handling open-ended, real-world VQA tasks, achieving high scores across dimensions like correctness, contextual understanding, and temporal understanding.

### v2

@@ Abstract
Recent advancements in multimodal large language models (MLLMs) have opened new avenues for video understanding. However, achieving high fidelity in zero-shot video tasks remains challenging. Traditional video processing methods rely heavily on fine-tuning to capture nuanced spatial-temporal details, which incurs significant data and computation costs. In contrast, training-free approaches, though efficient, often lack robustness in preserving context-rich features across complex video content. To this end, we propose DyTo, a novel dynamic token merging framework for zero-shot video understanding that adaptively optimizes token efficiency while preserving crucial scene details. DyTo integrates a hierarchical frame selection and a bipartite token merging strategy to dynamically cluster key frames and selectively compress token sequences, striking a balance between computational efficiency with semantic richness. Extensive experiments across multiple benchmarks demonstrate the effectiveness of DyTo, achieving superior performance compared to both fine-tuned and training-free methods and setting a new state-of-the-art for zero-shot video understanding. Code is available at https://github.com/Jam1ezhang/DYTO.
@@ 1 Introduction
Figure 1: Comparison with two SOTA training-free video-based LLM decoding methods over three different model backbones across five video benchmarks. DyTo and other baselines are marked using solid (—) and dashed (- - -) lines, respectively. DyTo outperforms existing training-free SOTA methods on almost all the benchmarks and achieves even better performance than most SFT-based methods.
@@ 2.1 Video Large Language Models
Video language model has witnessed significant advancements over the past years [20, 6, 46, 32, 38]. Video-ChatGPT [26] retrieves features from each frame and subsequently combines them using two operations for spatial and temporal pooling prior to feeding them into a large language model. LLaMA-VID [18] utilized a dual token approach to effectively compress the video token by differentiating between context and content. Vista-LLaMA [23] presented EDVT-Attention along with a sequential vision projector that emphasizes visual tokens while decreasing temporal tokens by sequentially merging them using a Q-former. Video-LLaVA [20] aligns the encoders for images and videos beforehand, enabling shared projections and joint training across both image and video tasks, thus mapping them into the language space. VideoChat [15] utilized cross-attention to condense video tokens alongside user inquiries and conversational context. Video-LLaMA [43] introduces a Video Q-Former and an Audio Q-Former, allowing for the integration of multiple modalities in video understanding. In contrast, Video-LLaMA2 [6] designs a spatial-temporal convolution connector to replace the Q-Former for spatial-temporal representation learning. Chat-UniVi [9] developed a unified model for images and videos that uses dynamic token merging with k-NN to simplify spatial and temporal tokens. SeViLA [41] focused on extracting key frames that are relevant to the inquiries and analyzed the video by transforming these keyframes into video tokens.
@@ 3 Method
Figure 2: The overview of DyTo, a training-free model built upon image-based MLLM without any fine-tuning. Specifically, DyTo first divides the video into $K$ clusters using the [CLS] token (pink block). Then the dynamic bipartite merging module samples frames from each cluster and controls the final output length as $Z$, resulting in better balance between computational efficiency and semantic richness.
@@ 3.1 Coarse-grained Hierarchical Clustering
The 1-NN temporal graph $G$ defines an adjacency matrix where each node is connected to its nearest neighbor based on the temporally weighted distances $W_{t}$. We make the links symmetric by setting $G(j,i)=1$ to encode space-temporal distances and form connected components to clusters conveniently. The connected components of Graph in Equ.2 automatically partition the data into discovered clusters. We then recursively merge these clusters according to the similarity of their feature averages. Throughout this merging process, we identify multiple clustering results and select the second-largest one as the final segmentation of the video. We denote by $n$ the index set of frames belonging to the $n_{th}$ cluster. Consequently, the set of frames in the $n_{th}$ cluster can be written as:
$V_{all}=\{V_{n}|n\in\{1,2,...,K\}\}$ (3)
We uniformly sample frames from every cluster and combine them as the keyframe sequence $V_{s}=\{v_{1},v_{2},...,v_{K}\}\in\mathbb{R}^{K\times(L-1)\times D}$ from the segmentations. The segmentation can yield the distribution of events contained in the current video sequence, thereby guiding the fine-grained merging in subsequent steps.
@@ 3.2 Fine-grained Dynamic Bipartite Merging
Previous method [11, 39, 35] either simply concatenates frames and designs prompts for the VLM or uses pooling method on each image token to construct the video representation. DyTo propose a novel method to effectively construct video representations from selected frames. After acquiring the segmentations of the frame sequence through the previously described method, we implement a dynamic bipartite token merging approach inspired by ToMe [2] to minimize the number of visual tokens. For each individual frame in $v_{k}$, we sequentially divide the $R_{i}$ tokens into two non-overlap tokens sets $\mathbb{P}$ with $r_{i}$ tokens and $\mathbb{Q}$ with $R_{i}-r_{i}$ tokens at the $i^{th}$ steps, where initial $R_{0}=L-1$(exclude CLS token). We set the $r_{i}=R_{0}-\frac{Z}{K}$ to dynamically merge the image to preserve more tokens under the fixed visual token length $Z$. To obtain the similarity scores, each visual token is split into $H$ heads along channel dimensions, each with $D/H$ channels. The similarity score for each token pair is obtained by averaging the cosine similarity scores over all heads following Eq. (4)
@@ 4 Experiment
Method LVLM Base Frame NExTQA EgoSchema IntentQA VideoMME MVBench
Size Model Length acc. (%) acc. (%) acc. (%) acc. (%) acc. (%)
Models fine-tuned with video data
SeViLA [40] 7B CLIP-L + FlanT5 8 63.6 25.7 60.9 44.6 -
Video-LLaVA [20] 7B ViT-L + Vicuna 8 - 38.4 - 40.4 41.0
LLaMA-VID [19] 7B EVA-G + Vicuna 1fps - 38.5 - - 41.9
InternVideo2 [34] 7B VideoMAEv2-g+InternVL 16 59.1 32.1 - 41.6 -
LLaVA-NeXT-Video [45] 7B CLIP-L + Vicuna 32 - 43.9 - 46.5 33.7
VideoLLaMA2 [6] 7B CLIP-L+ Mistral-Instruct 32 - 51.7 - 46.6 54.6
VideoChat2 [16] 7B UMT-L+Vicuna 16 61.7 - 59.0 39.5 51.1
LLaVA-OneVision [12] 7B SigLIP+Qwen-2 32 79.4 60.1 - 58.2 56.7
VideoLLaMA2 [6] 46.7B CLIP-L+ Mistral-Instruct 32 - 53.3 - 47.9 53.9
LLaVA-NeXT-Video [45] 32B CLIP-L + Qwen1.5 32 77.3 60.9 - 60.2 -
Training-free approaches
IG-VLM [11] 7B 6 63.1 35.8 60.1 39.8 41.3
SlowFast [39] 7B LLaVA-NeXT-image 50 64.2 45.5 60.1 40.4 44.8
DyTo 7B dynamic 65.7 48.6 61.6 41.2 45.2
IG-VLM [11] 8B 6 79.9 65.6 81.8 49.4 65.3
SlowFast [39] 8B InternVL2 50 77.2 60.7 79.8 48.0 57.0
DyTo 8B dynamic 81.4 67.6 83.0 59.3 66.2
IG-VLM [11] 9B 6 53.2 24.8 51.4 26.5 32.0
SlowFast [39] 9B Qwen-VL-Chat 50 48.4 17.6 45.1 22.1 31.7
DyTo 9B dynamic 53.4 26.0 51.7 27.4 34.3
IG-VLM [11] 26B 6 80.6 56.0 83.1 50.8 66.4
SlowFast [39] 26B InternVL2 50 79.2 54.8 82.8 49.4 61.8
DyTo 26B dynamic 81.1 59.2 83.6 53.0 68.1
IG-VLM [11] 34B 6 70.9 53.6 65.3 52.0 48.4
SlowFast [39] 34B LLaVA-NeXT-image 50 71.9 55.8 66.2 53.2 51.2
DyTo 34B dynamic 72.9 56.8 67.5 53.4 52.9
Table 1: Structured VQA benchmarks results comparing DyTo with SOTA training-free approaches as well as models that have been fine-tuned with additional video data from various LVLM architectures.
@@ 4.1 Evaluation Benchmarks
For structured VQA, we evaluate DyTo on a diverse set of multiple-choice benchmarks, including NExTQA [36], EgoSchema [28], IntentQA [13], VideoMME [7], and MVBench [16], all of which are designed to quantify model’s capability on video understanding by selecting the correct answer among predefined options. Collectively, these benchmarks provide a comprehensive evaluation of DyTo to interpret complex, multimodal data and to select accurate, contextually rich responses across varying levels of task structures. Notably, we perform experiments on the VideoMME [7] benchmark under the “w/o subs” configuration, which restricts access to subtitles, thereby isolating the model’s reliance on visual and temporal cues in video understanding.
@@
We also assess DyTo on the open-ended VQA tasks, specifically, we evaluate the zero-shot performance of DyTo on MSVD-QA [3], MSRVTT-QA [37], TGIF-QA [17], ANet-QA [42], and Video-ChatGPT Generation (VCG) [24] benchmarks. These benchmarks require the model to autonomously generate free-form responses, simulating real-world question-answering contexts that demand nuanced understanding of video content. Specifically, for VCG benchmark, we assess across five key dimensions, including Correctness of Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU), and Consistency (CO). Following Wu [35], we utilize GPT-3.5-Turbo-0125 to ensure fair comparisons with other methods.
@@ 4.2 Experimental Setup
In our approach, we uniformly sample each video to $N=100$ frames. Each frame is resized to match the input dimensions of different visual encoder, which then outputs visual tokens plus a [CLS] token. Following Sarfraz et al. [31], we collect the [CLS] token as the coarse-grained feature for each frame and subsequently partition the sequence into $K$ clusters, using recursive iterations of hierarchical clustering3.1) to group semantically similar frames.
Due to hardware limitations, we set the visual token sequence length $Z$ to either 3680 or 7200, corresponding to model sizes of 7B and 34B, respectively. To optimize feature representation and mitigate token redundancy, we dynamically adjust the merge ratio based on the formulation $r=Z/K$.
@@ 4.3 Main Results
As shown in Table 1, DyTo outperforms all the training-free approaches as well as all the fine-tuned models across all benchmarks by clear margins. Notably, DyTo demonstrated its adaptability to diverse video understanding contexts, achieving superior accuracy even compared to models that employ extensive fine-tuning. For example, on NExTQA [36], DyTo sets a new state-of-the-art performance by achieving 81.4% correctness when pairing with InternVL2-8B base model, which is significantly higher than the 59.1% accuracy achieved by InternVideo2 [34]. Similarly, performance on EgoSchema [28], IntentQA [13], VideoMME [7], and MVBench [16] illustrate the effectiveness of DyTo in handling task-specific reasoning within video content.
@@
Method LVLM Base Frame MSVD-QA MSRVTT-QA TGIF-QA ANet-QA Video-ChatGPT (VCG Benchmark)
Size Model Length acc./score acc./score acc./score acc./score CI DO CU TU CO Average
Models fine-tuned with video data
Video-ChatGPT[25] 7B CLIP-L+Vicuna 6 64.9/3.5 49.3/2.9 51.4/3.0 35.2/2.7 2.50 2.57 2.69 2.16 2.20 2.42
VideoGPT+[26] 3.8B CLIP-L+Phi-3-mini 16 72.4/3.9 60.6/3.6 74.6/4.1 50.6/3.6 3.27 3.18 3.74 2.83 3.39 3.28
Video-LLava[20] 7B ViT-L + Vicuna 8 70.7/3.9 59.2/3.5 70.0/4.0 45.3/3.3 - - - - - -
MovieChat[32] 7B CLIP-G+Vicuna 2048 75.2/3.8 52.7/2.6 - 45.7/3.4 2.76 2.93 3.01 2.24 2.42 2.67
LLama-VID[19] 13B EVA-G+Vicuna 1fps 69.7/3.7 57.7/3.2 - 47.4/3.3 2.96 3.00 3.53 2.46 2.51 2.89
VideoChat2[16] 7B UMT-L+Vicuna 16 70.0/3.9 54.1/3.3 - 49.1/3.3 3.02 2.88 3.51 2.26 2.81 2.98
Vista-LLAMA[23] 7B CLIP-L+Vicuna 16 65.3/3.6 60.5/3.3 - 48.3/3.3 2.44 2.64 3.18 2.26 2.31 2.57
Video-LLama2[6] 7B CLIP-L+Mistral-Instruct 32 70.9/3.8 - - 50.2/3.3 3.16 3.08 3.69 2.56 3.14 3.13
PLLaVA[38] 7B CLIP-L+Vicuna 16 76.6/4.1 62.0/3.5 77.5/4.1 56.3/3.5 - - - - - -
LLaVA-NeXT-Video[45] 7B CLIP-L + Vicuna 32 - - - 53.5/3.5 - - - - - -
Video-LLAMA2[6] 46.7B CLIP-L+Mistral-Instruct 32 70.5/3.8 - - 50.3/3.4 3.08 3.11 3.64 2.67 3.26 3.15
LLaVA-NeXT-Video[45] 34B CLIP-L + Qwen1.5 32 - - - 58.8/3.4 - - - - - -
Training-free approaches
IG-VLM-7B[11] 7B LLaVA-NeXT-image 6 78.3/3.9 63.7/3.4 72.7/4.0 53.8/3.2 3.11 2.78 3.48 2.44 3.29 3.03
IG-VLM-34B[11] 34B LLaVA-NeXT-image 6 79.6/4.1 62.4/3.5 79.1/4.2 58.4/3.5 3.21 2.87 3.54 2.51 3.34 3.09
SlowFast-7B[39] 7B LLaVA-NeXT-image 50 78.7/3.9 66.2/3.4 77.5/4.0 53.9/3.1 3.09 2.70 3.57 2.52 3.35 3.04
SlowFast-34B[39] 34B LLaVA-NeXT-image 50 78.7/4.1 67.1/3.7 80.6/4.3 58.8/3.5 3.48 2.96 3.84 2.70 3.54 3.30
DyTo 7B LLaVA-NeXT-iamge dynamic 77.6/3.9 64.1/3.4 78.0/4.0 54.3/3.2 3.12 2.93 3.72 2.52 3.52 3.16
DyTo 34B LLaVA-NeXT-image dynamic 79.6/4.1 66.2/3.6 80.7/4.2 59.0/3.5 3.45 2.95 3.82 2.72 3.63 3.32
@@
In open-ended settings, as shown in Table 2, DyTo also demonstrates competitive zero-shot capabilities. It consistently outperforms existing methods on benchmarks including MSVD-QA [3], MSRVTT-QA [37], TGIF-QA [17], and ANet-QA [42]. Notably, DyTo performs exceptionally well on Video-ChatGPT [24], surpassing even those methods that require fine-tuning, highlighting its robustness in zero-shot settings. DyTo ’s ability to maintain high accuracy without fine-tuning underscores the strength of its adaptive framework in handling open-ended, real-world VQA tasks, achieving high scores across dimensions like correctness, contextual understanding, and temporal understanding.